Here is the question as stated (this is not from a textbook, and I can't find an explanation for a problem like it in any textbooks I have):
"Write an equation in standard form for the line parallel to x - 2y + 7 = 0 and contains the point (-4,0)."
I am just unsure how to solve this problem. I have converted problems from slope intercept form to standard, and standard to slope intercept, but there hasn't been an extra component (like the +7 in this one.) I am also not sure how to do this when I only have one set of coordinates. I tried figuring out the value of x and the value of y by (if x=0 then y=?, and if y = 0 then x=?) but I have two problems. One is, the numbers I get (x= -7 and y= -7/2) don't make the problem true. And even if they did, I don't know how to use them to help me make the original equation a line parallel to another point. Help!

Answer by stanbon(48522)

Write an equation in standard form for the line parallel to x - 2y + 7 = 0 and contains the point (-4,0).
----------
The given equation can be written in the form y=(1/2)x+7/2
------------
so the equation you want has the same slope=1/2
The form is y=mx+b where y=0,x=-4 and m=1/2; solve for "b" as follows:
0=(1/2)(-4)+b
b=2
EQUATION:
y=(1/2)x+2
Now put it in standard form which is ax+by=c.
Multiply thru by 2 to get:
2y=x+4
x-2y=-4 is the equation in standard form.
==============
Cheers,
San H.